OpenStudy (anonymous):

is s^6+64 factored to be (s^2)(s^4-4s^4+16)

OpenStudy (raden):

first, setting s^6+64 as (s^2)^3+4^3

OpenStudy (raden):

then use the formula : a^3 + b^3 = (a+b)(a^2 - ab + b^2)

OpenStudy (anonymous):

okay

OpenStudy (raden):

in this case just change a = s^2 and b = 4 now distribute them into formula above

OpenStudy (raden):

did you get it ?